use the definition By definition of the Bernoulli numbers: noto prove that xcscx 2 (1) k-1 Bak (2²k 2) xc ²k (2k)!
Answer #1

We have:

sin z = 2 sin cos

sin x 2 sin - cos

\csc x=\dfrac 1 2 \csc \dfrac x 2 \sec \dfrac x 2=\dfrac 1 2 \csc \dfrac x 2 \dfrac 2 {e^{i \frac x 2} + e^{-i \frac x 2} }=\displaystyle \csc \dfrac x 2 \dfrac {e^{-i \frac x 2} } {1 + e^{-i x} }

csc x (1 +e-ir) = csc -e-ig

CSC.X= e-5-csc teir eir - e-ir 16 ix 2ix eir - 1 2 - 1 IF (ix) Bn (21x) B (2-n! - n!) 1 Bn (2 (ix) - (21x)) Bmixn-12 (1

Now, multiplying both sides by x and relabelling the index gives us:

8] - B2% (-1)*- * (22k I CSC T = 1 + \ (2)! 2)\square